OpenStudy (anonymous):

x^2+4x-8=o In simplest radical form

OpenStudy (anonymous):

b^2-4ac = 16+32 = 48 ==> x=-2+2sqrt(3) or =-2- 2sqrt(3)

OpenStudy (anonymous):

how did you get that ?